# A fast particle-field method for 3D Chemotaxis systemsJingyuan Hu, Zhongjian Wang, Jack Xin and Zhiwen Zhang

### v2

@@ 3.1 Particle-in-Cell Interpolation Framework
The space of trigonometric polynomials $\mathcal{T}_{H}$ is defined as:
$\mathcal{T}_{H}=\left\{f:f=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\right\}.$
The Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\hat{\beta}_{\mathbf{q}}$ defined in previous sections are specific evaluations:
(24) $\hat{\alpha}_{\mathbf{q}}=\mathcal{F}_{H}\hat{c}\left(\mathbf{q}\frac{2\pi}{L}\right),\quad\hat{\beta}_{\mathbf{q}}=\mathcal{F}_{H}\hat{\rho}\left(\mathbf{q}\frac{2\pi}{L}\right).$
The composition $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}$ forms a projection onto $\mathcal{T}_{H}$. For any function $f\in\mathcal{T}_{H}$, it holds that:
(25) $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}(f)=f.$
Additionally, for $\mathbf{q},\mathbf{q^{\prime}}\in U$ with $\mathbf{q}\neq\mathbf{q^{\prime}}$, the following properties are satisfied:
@@ 3.2 Computational Complexity and Algorithmic Details
Consider $P$ equally weighted particles at positions $\{\widehat{\mathbf{X}}_{p}\}_{p=1}^{P}\subset\Omega$ with total mass $M_{0}$. The empirical measure is shown as:
@@
With $H$ Fourier modes and $P$ particles, the complexity of particle-to grid discretization in Alg. 2 is $O(H^{3}\log H+P)$.
@@
For the inverse operation, given the values of a function $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$, we perform Alg. 3 to compute $I^{*}f(\mathbf{X}_{1}),\ldots,I^{*}f(\mathbf{X}_{P})$, where the interpolation operator $I^{*}$ is specified in Definition 3.2.
@@
The Grid-to-Particle transfer costs $O(P)$ (each particle interact with $|\Gamma|$ neighbors).
The computational complexity of the original method, dominated by all-to-all particle-frequency interactions scaling as $O(PH^{3})$, is reduced to $O(P+H^{3}\log H)$ by the PIC acceleration in Section 3.1. This involves local grid-projected density interpolation and gradient reconstruction, limiting particle interactions to neighboring grid points while handling frequency components globally. Periodic settings are applied to the boundary grids. To enhance numerical stability, we implement a spectral cutoff that removes high-frequency noise exceeding the threshold $H_{0}\cdot\frac{2\pi}{L}$.
Algorithm 4 SIPF-PIC
@@
The single-step complexity of Algorithm 4 is
$O\left(H^{3}\log H+P\right)$.
@@ 3.3 Bounds of Interpolation Error
Recall the empirical measure (27) and the discretization operator $I$ defined in Def. 3.1. We claim the following error bounds of $I$ under a Fourier transform.
@@
For any particle distribution $\{\mathbf{X}_{p}\}_{p=1}^{P}\subset\Omega$, the empirical density $\hat{\rho}^{(n-1)}$ is defined by (6). The frequencies $\mathbf{y}=\mathbf{q}\frac{2\pi}{L}$ for indices $\mathbf{q}\in U$. The interpolation errors satisfy:
(28) $|(\mathcal{F}_{H}I_{2}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{2}L^{-3}|\mathbf{y}|^{2}(\frac{L}{H})^{2},$
(29) $|(\mathcal{F}_{H}I_{4}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{4}L^{-3}|\mathbf{y}|^{4}(\frac{L}{H})^{4},$
where $C_{\eta}$ depends on interpolation order $\eta$.
@@
Let $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$. Then
(30) $\mathcal{F}_{H}\hat{\rho}(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}g_{\mathbf{y}}(\mathbf{X}_{p}),$
and
(31) $\mathcal{F}_{H}I(\hat{\rho})(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}\sum_{\gamma\in\Gamma}F_{p,\gamma}g_{\mathbf{y}}(\mathbf{X}_{p,\gamma}).$
Using a $d$-th order Taylor expansion with Lagrange remainder at $\mathbf{X}$, for multi-index $\mathbf{k}=(k_{1},k_{2},k_{3})$ and $D_{\mathbf{k}}=\partial^{|\mathbf{k}|}/\partial x_{1}^{k_{1}}\partial x_{2}^{k_{2}}\partial x_{3}^{k_{3}}$, we have:
@@
where $\xi$ lies between $\mathbf{X}$ and $\mathbf{x}$, and $\mathbf{k}!=k_{1}!k_{2}!k_{3}!$.
The exactness of $I_{\eta}$ for polynomials of degrees 0 to $\eta-1$ implies cancellation of lower-order terms, in other words,
(33) $\sum_{\gamma\in\Gamma}F_{\gamma}f(\mathbf{X}_{\gamma})=f(\mathbf{X})$
holds for all polynomials $f$ of degree at most $\eta-1$, yielding
(34) $g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}F_{\gamma}g_{\mathbf{y}}(\mathbf{X}_{\gamma})=\sum_{|\mathbf{k}|=\eta}\sum_{\gamma\in\Gamma}F_{\gamma}\frac{D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))}{\mathbf{k}!}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}.$
Since $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$, we have
(35) $|D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))|\leq|\mathbf{y}|^{\eta},$
for all $\mathbf{k},\mathbf{X},\xi,\gamma$.
For the case of $\eta=2$ and $\Gamma=\Gamma_{2}$, we have $F_{\gamma}\geq 0$, $\sum_{\gamma\in\Gamma}F_{\gamma}=1$ and
(36) $(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\leq\left(\frac{L}{H}\right)^{2}.$
Thus,
(37) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}F_{\gamma}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\sum_{|\mathbf{k}|=2}\frac{|\mathbf{y}|^{2}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{2}=\frac{9}{2}|\mathbf{y}|^{2}\left(\frac{L}{H}\right)^{2}.$
@@
We only need to handle the following:
(38) $\left|\sum_{\gamma\in\Gamma}F_{\gamma}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$ $\leq\left|\sum_{\gamma\in\Gamma_{2}}F_{\gamma}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|+\left|\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}F_{\gamma}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$
$\leq\left(\frac{L}{H}\right)^{4}\left(\sum_{\gamma\in\Gamma_{2}}F_{\gamma}-2^{4}\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}F_{\gamma}\right)$
$=\left(\frac{L}{H}\right)^{4}\left(1-17\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}F_{\gamma}\right).$
We partition the 24 outer grid points into 8 clusters based on their nearest inner vertices. For instance, the cluster containing $\mathbf{X}_{211},\mathbf{X}_{121},$ and $\mathbf{X}_{112}$ shares adjacency with the inner vertex $\mathbf{X}_{111}$. For each cluster, the coefficient combination satisfies:
(39) $(F_{211}+F_{121}+F_{112})=\lambda_{1}\lambda_{2}\lambda_{3}\left(\frac{3-\lambda_{1}^{2}-\lambda_{2}^{2}-\lambda_{3}^{2}}{6}\right)\leq\frac{1}{2}\lambda_{1}\lambda_{2}\lambda_{3}.$
Aggregating over all 8 clusters yields:
(40) $\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}F_{\gamma}\leq\frac{1}{2}\left(\lambda_{1}\lambda_{2}\lambda_{3}+\cdots+(1-\lambda_{1})(1-\lambda_{2})(1-\lambda_{3})\right)=\frac{1}{2}.$
Combining these results, we establish the fourth-order error bound:
(41) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}F_{\gamma}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\frac{19}{2}\sum_{|\mathbf{k}|=4}\frac{|\mathbf{y}|^{4}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{4}=\frac{513}{16}|\mathbf{y}|^{4}\left(\frac{L}{H}\right)^{4}.$
The overall spectral error consequently satisfies:
@@
where $\eta=2$ (with $C_{2}=\frac{9}{2}$) and $\eta=4$ (with $C_{4}=\frac{513}{16}$) correspond to second-order and fourth-order interpolation respectively.
@@
Assume $f\in\\ C^{2}(\Omega)$. Given the values of $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$. For each $p$ from 1 to $P$, the error of the second-order interpolation scheme at position $\mathbf{X}_{p}$ is
@@
further assume $f\in C^{4}(\Omega)$, the error of the fourth-order interpolation scheme satisfies
@@
where $K_{2}=\frac{9}{2}\max_{|\mathbf{k}|=2}\|D_{\mathbf{k}}f\|_{\infty}$ is related to the second derivative of $f$, and $K_{4}=\frac{513}{16}\max_{|\mathbf{k}|=4}\|D_{\mathbf{k}}f\|_{\infty}$ is related to the fourth derivative of $f$.
4 Convergence of SIPF-PIC
In this section, we present the convergence result of the proposed SIPF-PIC method (Algorithm 4), which is listed as Theorem 4.16. Specifically, we introduce the following Brownian coupling technique to construct an auxiliary random variable $\widetilde{\mathbf{X}}$ that bridges the ground truth random variable $\mathbf{X}$ and the empirical random variable $\widehat{\mathbf{X}}$ maintained in the algorithm.
@@ 4 Convergence of SIPF-PIC
$\widetilde{\mathbf{X}}^{(n)}$ denotes the discrete random variable of particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}$, where each particle is taken with a probability of $\frac{1}{P}$.
@@
Furthermore, the auxiliary concentration $\tilde{c}^{(n)}$ and its Fourier coefficients $\tilde{\alpha}^{(n)}$ are updated through Alg. 4:
@@
The semi-discrete implicit reference concentration $c^{(n)}$ and its Fourier coefficients $\alpha^{(n)}$ are updated through:
@@
where the corresponding random variable $\mathbf{X}^{(n)}$ is updated through an Euler-Maruyama semi-discretization over the time interval, and $\rho^{(n)}$ denotes its distribution:
@@
The probability space in the algorithm is generated as follows: the initial value $(\mathbf{X}_{p}^{(0)})_{1\leq p\leq P}$; the diffusion term in the $n$-th step $(\mathbf{W}_{p}^{(n)})_{1\leq p\leq P}$, where $1\leq n\leq N_{T}$. When discussing probabilities or using expectation notation $\mathbb{E}$, we always operate within this space. The updates of particles $\widehat{\mathbf{X}}_{p}^{(n)}$ and auxiliary particles $\widetilde{\mathbf{X}}_{p}^{(n)}$ depend on this shared probability space, which constitutes the essence of the Brownian coupling technique.
@@
The frequency error tensor $e_{\mathbf{q}}^{(n)}$ is defined as
(54) $e_{\mathbf{q}}^{(n)}=\hat{\alpha}_{\mathbf{q}}^{(n)}-\tilde{\alpha}_{\mathbf{q}}^{(n)},$
where the Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\tilde{\alpha}_{\mathbf{q}}$ are defined through (10) and (48) for indices $\mathbf{q}\in U=\left\{-\frac{H}{2},\ldots,\frac{H}{2}-1\right\}^{3}$.
@@
The solution $\rho,c$ of (1) is periodic, that is, for $q_{1},q_{2},q_{3}\in\mathbf{Z}$,
@@
The functions $\rho$ and $\nabla\rho$ are square-integrable, i.e.,
@@
The first to third spatial derivatives of $c$ and $\hat{c}$ are bounded, that is, for any spatial multi-index $\mathbf{k}$ such that $|\mathbf{k}|=\eta$ where $\eta=1,2,3$, it holds that
@@
4.1 Preliminary Estimates and Main Convergence Proof
We outline the structure of the argument. In Section 4.1, we first present a series of lemmas that establish bounds on individual components of the error dynamics, and then prove Theorem 4.16, which constitutes the main convergence result for Algorithm 4. In Section 4.2, we finally prove that under specific parameter combinations, the smoothness of $\hat{c}$ in Assumption 4.5 can be removed, which implies that our algorithm is numerically stable.

### v3

@@ 3.1 Particle-in-cell interpolation framework
The Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\hat{\beta}_{\mathbf{q}}$ defined in (5) and (9) are specific evaluations: $\hat{\alpha}_{\mathbf{q}}=\mathcal{F}_{H}\hat{c}\left(\mathbf{q}\frac{2\pi}{L}\right),\quad\hat{\beta}_{\mathbf{q}}=\mathcal{F}_{H}\hat{\rho}\left(\mathbf{q}\frac{2\pi}{L}\right).$ The composition $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}$ forms a projection onto the space of trigonometric polynomials $\mathcal{T}_{H}:=\left\{f:f=\sum_{\mathbf{q}\in U}\hat{\alpha}_{\mathbf{q}}e^{i\frac{2\pi}{L}\mathbf{q}\cdot\mathbf{x}}\right\}$. For any function $f\in\mathcal{T}_{H}$, it holds that $\mathcal{F}_{H}^{-1}\circ\mathcal{F}_{H}(f)=f.$ Additionally, for $\mathbf{q},\mathbf{q^{\prime}}\in U$ with $\mathbf{q}\neq\mathbf{q^{\prime}}$, the following properties are satisfied:
@@ 3.2 Computational complexity and algorithmic details
Consider $P$
equally weighted particles at positions $\{\widehat{\mathbf{X}}_{p}\}_{p=1}^{P}\subset\Omega$ with total mass $M_{0}$. The empirical measure is shown as:
@@
With $H$ Fourier modes and $P$ particles, the complexity of particle-to-grid discretization in Algorithm 2 is $O(H^{3}\log H+P)$.
@@
For the inverse operation, given the values of a function $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$, we perform Algorithm 3 to compute $I^{*}f(\mathbf{X}_{1}),\ldots,I^{*}f(\mathbf{X}_{P})$, where the interpolation operator $I^{*}$ is specified in Definition 3.2.
@@
The grid-to-particle transfer costs $O(P)$ (each particle interacts with $|\Gamma|$ neighbors).
The computational complexity of the original method, dominated by all-to-all particle-frequency interactions scaling as $O(PH^{3})$, is reduced to $O(P+H^{3}\log H)$ via the PIC acceleration presented in Section 3.1. This acceleration involves local grid-projected density interpolation and gradient reconstruction, limiting particle interactions to neighboring grid points while handling frequency components globally. Periodic boundary conditions are imposed on the boundary grids. To enhance numerical stability, we implement a spectral cutoff that filters out high-frequency modes exceeding the threshold $H_{0}\cdot\frac{2\pi}{L}$.
Algorithm 4 SIPF-PIC method
@@
The single-step complexity of Algorithm 4 is $O\left(H^{3}\log H+P\right)$.
@@ 3.3 Bounds of interpolation error
Recall the empirical measure (24) and the discretization operator $I$ defined in Definition 3.1. We claim the following error bounds for $I$ under a Fourier transform.
@@
For any particle distribution $\{\mathbf{X}_{p}\}_{p=1}^{P}\subset\Omega$, the empirical density $\hat{\rho}^{(n-1)}$ is defined by (6). let the frequencies be $\mathbf{y}=\mathbf{q}\frac{2\pi}{L}$ for indices $\mathbf{q}\in U$. The interpolation errors satisfy:
(25) $|(\mathcal{F}_{H}I_{2}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{2}L^{-3}|\mathbf{y}|^{2}\left(\frac{L}{H}\right)^{2},$
(26) $|(\mathcal{F}_{H}I_{4}(\hat{\rho})-\mathcal{F}_{H}\hat{\rho})(\mathbf{y})|$ $\leq M_{0}C_{4}L^{-3}|\mathbf{y}|^{4}\left(\frac{L}{H}\right)^{4},$
where $C_{\eta}$ depends on the interpolation order $\eta$ with $\eta=2,4$.
@@
Let $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$. Then we have the Fourier transform
(27) $\mathcal{F}_{H}\hat{\rho}(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}g_{\mathbf{y}}(\mathbf{X}_{p}),\quad\mathcal{F}_{H}I(\hat{\rho})(\mathbf{y})=\frac{M_{0}}{PL^{3}}\sum_{p=1}^{P}\sum_{\gamma\in\Gamma}F_{p,\gamma}g_{\mathbf{y}}(\mathbf{X}_{p,\gamma}).$
Using a $d$-th order Taylor expansion with Lagrange remainder at $\mathbf{X}$, for multi-index $\mathbf{k}=(k_{1},k_{2},k_{3})$ and $D_{\mathbf{k}}=\partial^{|\mathbf{k}|}/\partial x_{1}^{k_{1}}\partial x_{2}^{k_{2}}\partial x_{3}^{k_{3}}$, we have
@@
where $\xi$ lies between $\mathbf{X}$ and $\mathbf{x}$, and $\mathbf{k}!=k_{1}!k_{2}!k_{3}!$. Throughout this paper, the notation $|\mathbf{k}|=k_{1}+k_{2}+k_{3}$ means the order of a multi-index $\mathbf{k}$, which is different from the norm $|\mathbf{x}|$ for a vector $\mathbf{x}$ in Euclidean space.
The exactness of $I_{\eta}$ for polynomials of degrees 0 to $\eta-1$ implies cancellation of lower-order terms, in other words, $\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}f(\mathbf{X}_{\gamma})=f(\mathbf{X})$ holds for all polynomials $f$ of degree at most $\eta-1$, where $\overline{F_{\gamma}}:=F_{\gamma}(\boldsymbol{\lambda}(\mathbf{X}))$, yielding:
(29) $g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})=\sum_{|\mathbf{k}|=\eta}\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}\frac{D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))}{\mathbf{k}!}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}.$
Since $g_{\mathbf{y}}(\mathbf{x})=e^{-i\mathbf{y}\cdot\mathbf{x}}$, we have $|D_{\mathbf{k}}g_{\mathbf{y}}(\mathbf{X}+\xi(\mathbf{X},\gamma))|\leq|\mathbf{y}|^{\eta},$ for all $\mathbf{k},\mathbf{X},\xi,\gamma$.
For the case of $\eta=2$ and $\Gamma=\Gamma_{2}$, we have $\overline{F_{\gamma}}\geq 0$, $\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}=1$ and $(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\leq\left(\frac{L}{H}\right)^{2}.$ Thus, we establish the second-order error bound
(30) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\sum_{|\mathbf{k}|=2}\frac{|\mathbf{y}|^{2}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{2}=\frac{9}{2}|\mathbf{y}|^{2}\left(\frac{L}{H}\right)^{2}.$
@@
We only need to handle the following
(31) $\left|\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$ $\leq\left|\sum_{\gamma\in\Gamma_{2}}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|+\left|\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}(\mathbf{X}_{\gamma}-\mathbf{X})^{\mathbf{k}}\right|$
$\leq\left(\frac{L}{H}\right)^{4}\left(\sum_{\gamma\in\Gamma_{2}}\overline{F_{\gamma}}-2^{4}\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\right)$
$=\left(\frac{L}{H}\right)^{4}\left(1-17\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\right).$
We partition the 24 outer grid points into 8 clusters based on their nearest inner vertices. For instance, the cluster containing $\mathbf{X}_{211},\mathbf{X}_{121},$ and $\mathbf{X}_{112}$ shares adjacency with the inner vertex $\mathbf{X}_{111}$. For each cluster, the coefficient combination satisfies
(32) $(\overline{F}_{211}+\overline{F}_{121}+\overline{F}_{112})=\lambda_{1}\lambda_{2}\lambda_{3}\left(\frac{3-\lambda_{1}^{2}-\lambda_{2}^{2}-\lambda_{3}^{2}}{6}\right)\leq\frac{1}{2}\lambda_{1}\lambda_{2}\lambda_{3}.$
Aggregating over all 8 clusters yields
(33) $\sum_{\gamma\in\Gamma\setminus\Gamma_{2}}\overline{F_{\gamma}}\leq\frac{1}{2}\left(\lambda_{1}\lambda_{2}\lambda_{3}+\cdots+(1-\lambda_{1})(1-\lambda_{2})(1-\lambda_{3})\right)=\frac{1}{2}.$
Combining these results, we establish the fourth-order error bound
(34) $\left|-g_{\mathbf{y}}(\mathbf{X})+\sum_{\gamma\in\Gamma}\overline{F_{\gamma}}g_{\mathbf{y}}(\mathbf{X}_{\gamma})\right|\leq\frac{19}{2}\sum_{|\mathbf{k}|=4}\frac{|\mathbf{y}|^{4}}{\mathbf{k}!}\left(\frac{L}{H}\right)^{4}=\frac{513}{16}|\mathbf{y}|^{4}\left(\frac{L}{H}\right)^{4}.$
The overall spectral error consequently satisfies
@@
where $\eta=2$ (with $C_{2}=\frac{9}{2}$) and $\eta=4$ (with $C_{4}=\frac{513}{16}$) correspond to second-order and fourth-order interpolation, respectively.
@@
Assume $f\in\\ C^{2}(\Omega)$. Given the values of $f$ on the grid $f\left(U\cdot\frac{L}{H}\right)$, for each p=1,,Pp=1,...,P, the error of the second-order interpolation scheme at position $\mathbf{X}_{p}$ satisfies
@@
if we further assume $f\in C^{4}(\Omega)$, the error of the fourth-order interpolation scheme satisfies
@@
where $K_{2}=\frac{9}{2}\max_{|\mathbf{k}|=2}\|D_{\mathbf{k}}f\|_{\infty}$ is associated with the second derivative of $f$, and $K_{4}=\frac{513}{16}\max_{|\mathbf{k}|=4}\|D_{\mathbf{k}}f\|_{\infty}$ is associated with the fourth derivative of $f$.
4 Convergence of the SIPF-PIC method
In this section, we present the convergence result of the proposed SIPF-PIC method (Algorithm 4), which is stated as Theorem 4.16. Specifically, we introduce the following Brownian coupling technique to construct an auxiliary random variable $\widetilde{\mathbf{X}}$ that bridges the ground truth random variable $\mathbf{X}$ and the empirical random variable $\widehat{\mathbf{X}}$ maintained in the algorithm.
@@ 4 Convergence of the SIPF-PIC method
$\widetilde{\mathbf{X}}^{(n)}$ denotes the discrete random variable consisting of particles $\widetilde{\mathbf{X}}_{1}^{(n)},\ldots,\widetilde{\mathbf{X}}_{P}^{(n)}$, where each particle is sampled with probability $\frac{1}{P}$.
@@
Furthermore, the auxiliary concentration $\tilde{c}^{(n)}$ and its Fourier coefficients $\tilde{\alpha}^{(n)}$ are updated via Algorithm 4:
@@
The semi-discrete implicit reference concentration $c^{(n)}$ and its Fourier coefficients $\alpha^{(n)}$ are updated as follows:
@@
where the corresponding random variable $\mathbf{X}^{(n)}$ is updated via an Euler-Maruyama semi-discretization over the time interval, and $\rho^{(n)}$ denotes its distribution
@@
The probability space in the algorithm is generated as follows: the initial values $(\mathbf{X}_{p}^{(0)})_{1\leq p\leq P}$, and the diffusion terms $(\mathbf{W}_{p}^{(n)})_{1\leq p\leq P}$ in the $n$-th step for $1\leq n\leq N_{T}$. When discussing probabilities or using the expectation notation $\mathbb{E}$, we always operate within this space. The updates of particles $\widehat{\mathbf{X}}_{p}^{(n)}$ and auxiliary particles $\widetilde{\mathbf{X}}_{p}^{(n)}$ depend on this shared probability space, which constitutes the essence of the Brownian coupling technique.
@@
The frequency error $e_{\mathbf{q}}^{(n)}$ is defined as $e_{\mathbf{q}}^{(n)}:=\hat{\alpha}_{\mathbf{q}}^{(n)}-\tilde{\alpha}_{\mathbf{q}}^{(n)}$, where the Fourier coefficients $\hat{\alpha}_{\mathbf{q}}$ and $\tilde{\alpha}_{\mathbf{q}}$ are defined via (10) and (41) for indices $\mathbf{q}\in U=\left\{-\frac{H}{2},\ldots,\frac{H}{2}-1\right\}^{3}$.
@@
The solution $\rho,c$ to (1) is periodic, that is, for $q_{1},q_{2},q_{3}\in\mathbf{Z}$,
@@
The function $\rho$ and its gradient $\nabla\rho$ are square-integrable, i.e.,
@@
The first to third order spatial derivatives of $c$ and $\hat{c}$ are bounded, i.e., for any spatial multi-index $\mathbf{k}$ such that $|\mathbf{k}|=\eta$ where $\eta=1,2,3$, it holds that
@@
4.1 Main convergence results
We outline the structure of the argument. In this section, we first present a series of lemmas that establish bounds for individual components of the error dynamics, and then prove Theorem 4.16, which constitutes the main convergence result of Algorithm 4. Then, in Section 4.2, we further prove that under specific parameter combinations, the smoothness assumption on $\hat{c}$ in Assumption 4.5 can be removed.